# ReCalKV: Low-Rank KV Cache Compression via Head Reordering and Offline Calibration

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable performance, yet their capability on long-context reasoning is often constrained by the excessive memory required to store the Key-Value (KV) cache. This makes KV cache compression an essential step toward enabling efficient long-context reasoning. Recent methods have explored reducing the hidden dimensions of the KV cache, but many introduce additional computation through projection layers or suffer from significant performance degradation under high compression ratios. To address these challenges, we propose ReCalKV, a post-training KV cache compression method that reduces the hidden dimensions of the KV cache. We develop distinct compression strategies for Keys and Values based on their different roles and varying importance in the attention mechanism. For Keys, we propose Head-wise Similarity–aware Reordering (HSR), which clusters similar heads and applies grouped SVD to the key projection matrix, reducing additional computation while preserving accuracy. For Values, we propose Offline Calibration and Matrix Fusion (OCMF) to preserve accuracy without extra computational overhead. Experiments show that ReCalKV outperforms existing low-rank compression methods, achieving high compression ratios with minimal performance loss. The code and models will be available at: https://github.com/XIANGLONGYAN/ReCalKV.
@@ 1 Introduction
Large language models (LLMs) vaswani2017attention ; llama1 ; llama3 have demonstrated outstanding performance across a wide range of tasks. To accelerate inference, modern LLMs cache intermediate Key-Value (KV) states, avoiding redundant computation during autoregressive decoding. However, as the input context length increases, the KV cache grows rapidly, leading to substantial memory overhead and bandwidth pressure. In practice, the KV cache often becomes the primary bottleneck for long-context inference. Consequently, compressing the KV cache is crucial for efficient and scalable deployment of LLMs.
@@
To reduce the size of the KV cache, recent works explore compression along multiple axes. Multi-query attention mqa and grouped-query attention ainslie2023gqa reduce the number of heads by sharing keys and values. Quantization methods atom lower KV cache’s precision, with some kivi ; kvquant pushing KV representations down to 2 bits. Others h2o ; snapkv ; xiao2023efficient ; dong2024get reduce the number of cached tokens by selecting only important ones, often based on attention scores. A few methods chang2025xkv ; liu2024minicache further compress across layers by reusing KV states. These approaches highlight the multi-dimensional nature of KV cache compression, targeting heads, precision, sequence length, and depth.
@@
Another line of work chang2024palu ; liu2024deepseek ; lin2024matryoshkakv explores KV cache compression from a different angle—by reducing the dimensionality of the hidden vector space of Keys and Values themselves. For example, MLA liu2024deepseek reduces memory via low-rank representations but requires training the model from scratch. Other approaches, such as EigenAttention saxena2024eigen and MastryohakaKV lin2024matryoshkakv , compress KV entries via projection into lower-dimensional subspaces. Although effective, their projection and reconstruction introduce extra decoding overhead, limiting applicability in latency-sensitive scenarios. Palu chang2024palu and LoRC zhang2024lorc address this issue by directly applying Singular Value Decomposition (SVD) to the KV projection layers, effectively compressing the hidden dimensions of the KV cache. This approach substantially reduces the overhead of runtime projection and reconstruction. However, both methods overlook the inherent asymmetry between Keys and Values in the attention mechanism, and their performance degrades notably under high KV cache compression ratios.
@@
To better explore KV cache compression along the hidden dimension, we conduct detailed analyses of the roles of Keys and Values in the attention mechanism. Our analyses reveal that: (i) Most modern LLMs llama1 ; llama2 ; llama3 use positional encoding—typically RoPE rope_roformer —which is applied to Keys. As a result, low-rank compressed Keys must be fully reconstructed during inference to enable positional encoding, which introduces additional computational overhead. This makes it essential to consider both accuracy and computational cost when compressing Key cache. (ii) We measure the Fisher information of the Key and Value projection layers. Our Fisher information analysis reveals that the Value projection matrices carry significantly higher importance than their Key counterparts, highlighting the crucial role of Value representations in the overall model behavior. Therefore, when compressing the Value cache, it is crucial to minimize accuracy degradation.
@@
Based on our analyses, we develop distinct compression strategies for Keys and Values based on their different roles and varying importance in the attention mechanism. For Keys, we propose Head-wise Similarity–aware Reordering (HSR). It first reorders attention heads based on their representation similarity, then groups similar heads together, and applies grouped SVD within each group. This reduces the Key cache size with low reconstruction overhead. Grouping similar heads helps the SVD better capture shared subspace structures, lowering approximation error and preserving accuracy. For Values, we propose Offline Calibration and Matrix Fusion (OCMF). We first apply SVD to the Value projection matrix, then calibrate the decomposed components using a small calibration dataset. Finally, we fuse the right factor of the SVD decomposition with the subsequent output projection matrix, eliminating the need for explicit reconstruction during inference.
@@
Extensive experiments demonstrate that ReCalKV consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing low-rank compression methods under various evaluation settings. For example, on the LLaMA-2-7B model llama2 evaluated on six zero-shot QA datasets, ReCalKV achieves an average accuracy of 63.64% under a 50% KV cache compression ratio, compared to 64.99% for the full-precision model—corresponding to only a 2% relative drop. Notably, since ReCalKV is orthogonal to quantization techniques, it can be seamlessly integrated with them to achieve even higher overall compression ratios.
@@
  • We propose the Offline Calibration and Matrix Fusion (OCMF) strategy for Value compression, preserving accuracy without added computation.
@@ 2 Related work
SVD-Based LLM Compression. Singular Value Decomposition (SVD) has been widely adopted for compressing LLMs by approximating weight matrices with low-rank factors. However, standard SVD NaiveSVD has been observed to cause notable performance degradation in practice. To address this, FWSVDfwsvd incorporates Fisher information to guide decomposition, while ASVDyuan2023asvd rescales weights to mitigate the impact of activation outliers. SVD-LLMsvdllm establishes a direct connection between singular values and compression loss, employing data whitening and selective truncation to minimize degradation. AdaSVDli2025adasvd further improves compression by introducing adaptive rank allocation and compensation mechanisms. Unlike prior methods that target model weights, we apply SVD to the key and value projections to compress the KV cache directly.
KV Cache Compression. To support long-context inference, various methods have been proposed to compress the KV cache. Quantization-based approaches are widely adopted: Atomatom performs per-token quantization, while WKVQuantwkvquant uses a two-level scheme for better accuracy. KIVIkivi and KVQuantkvquant combine per-token and per-channel quantization, with KVQuant further leveraging non-uniform and sparse techniques to handle outliers. In parallel, token eviction methods h2o ; snapkv ; xiao2023efficient ; dong2024get reduce memory by discarding less relevant tokens or retrieving only subsets during decoding. Beyond these strategies, several methods aim to reduce the hidden dimension of KV representations. DeepSeekV2liu2024deepseek uses MLA for built-in dimension reduction, but requires training from scratch. MatryoshkaKVlin2024matryoshkakv and Eigen-Attentionsaxena2024eigen project the KV cache into a low-rank space via additional projection matrices, at the cost of increased computation. HeadKV fu2024not uses SVD to reduce the number of KV heads, leading to improved efficiency. LoRCzhang2024lorc and Paluchang2024palu directly apply SVD to the KV projection matrices, reducing dimensionality with minimal architectural changes. Our method is a post-training low-rank KV compression approach and is orthogonal to quantization and token eviction, enabling easy integration for further compression gains.
Figure 1: Overview of the ReCalKV framework. The method consists of three key components: Head-wise Similarity-aware Reordering (HSR) for compressing Keys via grouped SVD, and Offline Calibration and Matrix Fusion (OCMF) for compressing Values without additional runtime overhead.
@@ 3.1 Preliminary
Singular Value Decomposition. Singular Value Decomposition svd is a classical matrix factorization technique widely used for low-rank approximation. Given a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD factorizes it into three components: $\mathbf{W}=\mathbf{U}{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, and $\boldsymbol{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix with non-negative singular values. To obtain a low-rank approximation of a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we apply singular value decomposition (SVD) and retain only the top $r$ singular values and their associated singular vectors. This results in an approximate factorization:
$\mathbf{W}\approx\mathbf{L}\mathbf{R},\quad\text{where}\quad\mathbf{L}=\mathbf{U}_{r}\boldsymbol{\Sigma}_{r}^{1/2},\quad\mathbf{R}=\boldsymbol{\Sigma}_{r}^{1/2}\mathbf{V}_{r}^{\top}.$ (1)
Here, $\mathbf{U}_{r}\in\mathbb{R}^{m\times r}$ and $\mathbf{V}_{r}\in\mathbb{R}^{n\times r}$ are the top-$r$ singular vectors of $\mathbf{W}$, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains the corresponding singular values. This yields two smaller matrices, $\mathbf{L}$ and $\mathbf{R}$, from the low-rank decomposition of $\mathbf{W}$. Given an input $\mathbf{x}\in\mathbb{R}^{1\times m}$, we compute the intermediate representation $\mathbf{z}=\mathbf{x}\mathbf{L}$ and store $\mathbf{z}\in\mathbb{R}^{1\times r}$ in the KV cache instead of the full output. During attention, the output is approximated by reconstructing $\mathbf{x}\mathbf{W}\approx\mathbf{z}\mathbf{R}$,. This approach reduces the KV cache size with a compression ratio of $r/n$, while maintaining a close approximation of the original computation.
Centered Kernel Alignment Similarity. Centered Kernel Alignment (CKA) kornblith2019similarity is a widely used metric for quantifying the similarity between two sets of representations. Given two matrices $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, CKA is computed by first forming their Gram (kernel) matrices $\mathbf{G_{X}}=\mathbf{X}\mathbf{X}^{\top}$ and $\mathbf{G_{Y}}=\mathbf{Y}\mathbf{Y}^{\top}$. These kernel matrices are then centered as follows:
@@
where $\text{HSIC}(\mathbf{X},\mathbf{Y})=\text{Tr}(\mathbf{\tilde{G_{X}}}\mathbf{\tilde{G_{Y}}})$ denotes the Hilbert-Schmidt Independence Criterion (HSIC). CKA ranges from 0 to 1, with higher values indicating greater similarity between the two matrics.
@@ 3.2 Head-wise Similarity-aware Reordering
Group-head Low-rank Decomposition. Following the grouped decomposition strategy proposed in Palu chang2024palu , we organize multiple attention heads into groups prior to performing SVD. Given a Key projection matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, where $n=h\cdot d_{h}$ corresponds to $h$ attention heads each of hidden dimension $d_{h}$, we divide $\mathbf{W}$ column-wise into $h$ submatrices, each representing a single head. Then we group every $s$ heads into one group, resulting in $g=h/s$ groups in total. For each group $j$, we construct a concatenated projection matrix $\mathbf{W}_{g_{j}}=\bigl{[}\mathbf{W}_{j,1},\dots,\mathbf{W}_{j,s}\bigr{]}$, where each $\mathbf{W}_{j,k}\in\mathbb{R}^{m\times d_{h}}$ is the projection matrix of the $k$-th head in group $j$, and thus $\mathbf{W}_{g_{j}}\in\mathbb{R}^{m\times(s\cdot d_{h})}$. Instead of applying SVD to the entire projection matrix at once, we apply low-rank approximation to the grouped matrix: $\mathbf{W}_{g_{j}}\approx\mathbf{L}_{g_{j}}\mathbf{R}_{g_{j}}$, where $\mathbf{L}_{g_{j}}\in\mathbb{R}^{d\times r_{g}}$ and $\mathbf{R}_{g_{j}}\in\mathbb{R}^{r_{g}\times(d_{h}\cdot s)}$. During inference, the latent representation shared across all heads in the group is computed as: $\mathbf{z}_{g_{j}}=\mathbf{x}\mathbf{L}_{g_{j}}$, and the projected outputs for individual heads are reconstructed via:
$[\mathbf{y}_{j,1},\dots,\mathbf{y}_{j,s}]=\mathbf{z}_{g_{j}}\mathbf{R}_{g_{j}}.$ (4)
@@
CKA-based Head Similarity. A key question in grouped SVD is how to assign attention heads into groups so as to minimize the reconstruction error. Empirically, we observe that grouping heads with similar left singular subspaces results in lower approximation error, as these heads tend to share more common representational components and thus benefit from joint compression. To quantify head similarity, we adopt centered kernel alignment (CKA) kornblith2019similarity , a robust and widely used metric for comparing representation subspaces. Specifically, we compute the pairwise CKA similarity between all attention heads, yielding a symmetric similarity matrix $\mathbf{S}\in\mathbb{R}^{h\times h}$, where $h$ is the total number of heads:
@@
Head Reordering. Based on the similarity matrix $\mathbf{S}$, we perform head grouping by prioritizing pairs with high mutual similarity. Specifically, we adopt a greedy strategy that iteratively selects the head pair with the highest CKA similarity and assigns them to the same group, subject to a fixed group size constraint (e.g., 4 heads per group when $h=32$). Remaining unassigned heads are then added to existing groups with available capacity, ensuring that all heads are eventually grouped. This head reordering process encourages heads with similar representational structure to share SVD decompositions, effectively reducing approximation error during grouped compression. As shown in Figure 2, we visualize the CKA similarity matrices before and after head reordering. It can be observed that, after reordering, heads assigned to adjacent positions exhibit higher mutual similarity, indicating that similar heads are more effectively grouped together. By aligning structurally coherent heads within each group, we improve both the compactness and accuracy of the resulting low-rank representation. To ensure decoding equivalence, we apply an inverse reordering to restore the original head order after grouped computation (see Figure 3).
3.3 Offline Calibration and Matrix Fusion

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance, but their long-context reasoning remains constrained by the excessive memory required for the Key-Value (KV) cache. This makes KV cache compression a critical step toward efficient long-context inference. Recent methods have explored low-rank techniques to reduce the hidden size of the KV cache. However, they neglect the distinct roles and varying importance of Keys and Values, leading to significant performance drops under high compression. To address this, we propose ReCalKV, a post-training low-rank KV cache compression approach with tailored strategies for Keys and Values. For Keys, we propose Head-wise Similarity–aware Reordering (HSR), which clusters structurally similar heads into groups, enabling more accurate low-rank approximation via grouped SVD. For Values, we propose Offline Value Calibration (OVC), which efficiently calibrates the value projection matrix using calibration data without training, ensuring an accurate representation of contextual information. Extensive experiments show that ReCalKV consistently outperforms existing low-rank compression methods, achieving high compression ratios with minimal performance loss. The code and models will be available at: https://github.com/XIANGLONGYAN/ReCalKV.
@@ 1 Introduction
Large language models (LLMs) (Vaswani, 2017; Touvron et al., 2023a; Dubey et al., 2024) have demonstrated outstanding performance across a wide range of tasks. To accelerate inference, modern LLMs cache intermediate Key-Value (KV) states, avoiding redundant computation during autoregressive decoding. However, as the input context length increases, the KV cache grows rapidly, leading to substantial memory overhead and bandwidth pressure. In practice, the KV cache often becomes the primary bottleneck for long-context inference. Consequently, compressing the KV cache becomes essential for enabling efficient and scalable deployment of LLMs across real-world applications.
@@
To reduce the size of the KV cache, recent works explore compression along multiple axes. Multi-query attention (Shazeer, 2019) and grouped-query attention (Ainslie et al., 2023) reduce the number of heads by sharing keys and values. Quantization methods (Zhao et al., 2023) lower KV cache’s precision, with some (Liu et al., 2024c; Hooper et al., 2024) pushing KV representations down to 2 bits. Others (Zhang et al., 2023; Li et al., 2024; Xiao et al., 2023; Dong et al., 2024) reduce the number of cached tokens by selecting only important ones, often based on attention scores. A few methods (Chang et al., 2025; Liu et al., 2024b) further compress across layers by reusing KV states. These methods reveal the multi-dimensional structure of KV cache compression.
@@
Another line of work (Chang et al., 2024; Liu et al., 2024a; Lin et al., 2024) explores KV cache compression from a different angle—by reducing the dimensionality of the hidden vector space of Keys and Values themselves. For example, MLA (Liu et al., 2024a) reduces memory via low-rank representations but requires training the model from scratch. Other approaches, such as EigenAttention (Saxena et al., 2024) and MastryohakaKV (Lin et al., 2024), compress KV entries via projection into lower-dimensional subspaces. Although effective, their projection and reconstruction introduce extra decoding overhead, limiting applicability in latency-sensitive scenarios. Palu (Chang et al., 2024) and LoRC (Zhang et al., 2024) address this issue by directly applying Singular Value Decomposition (SVD) to the KV projection layers, effectively compressing the hidden dimensions of the KV cache. This approach substantially reduces the overhead of runtime projection and reconstruction. However, both methods overlook the inherent asymmetry between Keys and Values in the attention mechanism, and their performance degrades notably under high KV cache compression ratios.
@@
To better explore KV cache compression along the hidden dimension, we conduct detailed analyses of the roles of Keys and Values in the attention mechanism. Our analyses reveal that: (i) Most modern LLMs (Touvron et al., 2023a; b; Dubey et al., 2024) use positional encoding, typically RoPE (Su et al., 2024), which is applied to Keys. As a result, low-rank compressed Keys must be fully reconstructed during inference to enable positional encoding, which introduces additional computational overhead. This makes it essential to consider both accuracy and computational cost when compressing Key cache. (ii) We measure the Fisher information of the Key and Value projection layers. Our Fisher information analysis reveals that the Value projection matrices carry significantly higher importance than their Key counterparts, highlighting the crucial role of Value representations in the overall model behavior. Therefore, minimizing accuracy degradation is critical when compressing the Value cache.
@@
Based on our analyses, we develop distinct compression strategies for Keys and Values based on their different roles and varying importance in the attention mechanism. For Keys, we propose Head-wise Similarity–aware Reordering (HSR). It first reorders attention heads based on their representation similarity, then groups similar heads together, and applies grouped SVD within each group. This reduces the Key cache size with low reconstruction overhead. Grouping similar heads helps the SVD better capture shared subspace structures, lowering approximation error and preserving accuracy. For Values, we propose Offline Calibration Value (OVC). We first apply SVD to the Value projection matrix and then calibrate the decomposed components with a small calibration dataset to preserve Value accuracy. We also fuse the right factor of the SVD decomposition into the subsequent output projection matrix, removing the need for explicit reconstruction during inference.
@@
Extensive experiments demonstrate that ReCalKV consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing low-rank compression methods under various evaluation settings. For example, on the LLaMA-2-7B model (Touvron et al., 2023b) evaluated on six zero-shot QA datasets, ReCalKV achieves an average accuracy of 63.64% under a 50% KV cache compression ratio, compared to 64.99% for the full-precision model—corresponding to only a 2% relative drop. Notably, since ReCalKV is orthogonal to quantization techniques, it can be seamlessly integrated with them to achieve even higher overall compression ratios.
@@
  • We propose an Offline Value Calibration (OVC) strategy for compressing the Value cache, preserving accuracy without introducing additional inference overhead.
@@ 2 Related work
SVD-Based LLM Compression. Singular Value Decomposition (SVD) has been widely adopted for compressing LLMs by approximating weight matrices with low-rank factors. However, standard SVD (Noach & Goldberg, 2020) has been observed to cause notable performance degradation in practice. To address this, FWSVD (Hsu et al., 2022) incorporates Fisher information to guide decomposition, while ASVD (Yuan et al., 2023) rescales weights to mitigate the impact of activation outliers. SVD-LLM (Wang et al., 2024) establishes a direct connection between singular values and compression loss, employing data whitening and selective truncation to minimize degradation. AdaSVD (Li et al., 2025) further improves compression by introducing adaptive rank allocation and compensation mechanisms. Unlike prior methods that target model weights, we apply SVD to the key and value projections to compress the KV cache directly.
KV Cache Compression. To support long-context inference, various methods have been proposed to compress the KV cache. Quantization-based approaches are widely adopted: Atom (Zhao et al., 2023) performs per-token quantization, while WKVQuant (Yue et al., 2024) uses a two-level scheme for better accuracy. KIVILiu et al. (2024c) and KVQuant (Hooper et al., 2024) combine per-token and per-channel quantization, with KVQuant further leveraging non-uniform and sparse techniques to handle outliers. In parallel, token eviction methods (Zhang et al., 2023; Li et al., 2024; Xiao et al., 2023; Dong et al., 2024) reduce memory by discarding less relevant tokens or retrieving only subsets during decoding. Beyond these strategies, several methods aim to reduce the hidden dimension of KV representations. DeepSeek V2 (Liu et al., 2024a) uses MLA for built-in dimension reduction, but requires training from scratch. MatryoshkaKV (Lin et al., 2024) and Eigen-Attention (Saxena et al., 2024) project the KV cache into a low-rank space via additional projection matrices, at the cost of increased computation. HeadKV (Fu et al., 2024) uses SVD to reduce the number of KV heads, leading to improved efficiency. LoRC (Zhang et al., 2024) and Palu (Chang et al., 2024) directly apply SVD to the KV projection matrices, reducing dimensionality with minimal architectural changes. Our method is a post-training low-rank KV compression approach and is orthogonal to quantization and token eviction, enabling easy integration for further compression gains.
Figure 1: Overview of the ReCalKV framework. The method consists of three key components: Head-wise Similarity-aware Reordering (HSR) for compressing Keys via grouped SVD, and Offline Value Calibration (OVC) for compressing Values without additional runtime overhead.
@@ 3.1 Preliminary
Singular Value Decomposition. Singular Value Decomposition (Golub et al., 1987) is a classical matrix factorization technique widely used for low-rank approximation. Given a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD factorizes it into three components: $\mathbf{W}=\mathbf{U}{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, and $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix with non-negative singular values. To obtain a low-rank approximation of a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we apply SVD and retain only the top $r$ singular values and their associated singular vectors. This results in an approximate factorization:
$\mathbf{W}\approx\mathbf{L}\mathbf{R},\quad\text{where}\quad\mathbf{L}=\mathbf{U}_{r}\bm{\Sigma}_{r}^{1/2},\quad\mathbf{R}=\bm{\Sigma}_{r}^{1/2}\mathbf{V}_{r}^{\top}.$ (1)
Here, $\mathbf{U}_{r}\in\mathbb{R}^{m\times r}$ and $\mathbf{V}_{r}\in\mathbb{R}^{n\times r}$ are the top-$r$ singular vectors of $\mathbf{W}$, and $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains the corresponding singular values. This yields two smaller matrices, $\mathbf{L}$ and $\mathbf{R}$, from the low-rank decomposition of $\mathbf{W}$. Given an input $\mathbf{x}\in\mathbb{R}^{1\times m}$, we compute the intermediate representation $\mathbf{z}=\mathbf{x}\mathbf{L}$ and store $\mathbf{z}\in\mathbb{R}^{1\times r}$ in the KV cache instead of the full output. During attention, the output is approximated by reconstructing $\mathbf{x}\mathbf{W}\approx\mathbf{z}\mathbf{R}$. This approach reduces the KV cache size with a compression ratio of $r/n$, while maintaining a close approximation of the original computation.
Centered Kernel Alignment Similarity. Centered Kernel Alignment (CKA) (Kornblith et al., 2019) is a widely used metric for quantifying the similarity between two sets of representations. Given two matrices $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, CKA is computed by first forming their Gram (kernel) matrices $\mathbf{G_{X}}=\mathbf{X}\mathbf{X}^{\top}$ and $\mathbf{G_{Y}}=\mathbf{Y}\mathbf{Y}^{\top}$. These kernel matrices are then centered as follows:
@@
[…]
@@ 3.2 Head-wise Similarity-aware Reordering
Group-head Low-rank Decomposition. Following the grouped decomposition strategy proposed in Palu (Chang et al., 2024), we organize multiple attention heads into groups prior to performing SVD. Given a Key projection matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, where $n=h\cdot d_{h}$ corresponds to $h$ attention heads each of hidden dimension $d_{h}$, we divide $\mathbf{W}$ column-wise into $h$ submatrices, each representing a single head. Then we group every $s$ heads into one group, resulting in $g=h/s$ groups in total. For each group $j$, we construct a concatenated projection matrix $\mathbf{W}_{g_{j}}=\bigl[\mathbf{W}_{j,1},\dots,\mathbf{W}_{j,s}\bigr]$, where each $\mathbf{W}_{j,k}\in\mathbb{R}^{m\times d_{h}}$ is the projection matrix of the $k$-th head in group $j$, and thus $\mathbf{W}_{g_{j}}\in\mathbb{R}^{m\times(s\cdot d_{h})}$. Instead of applying SVD to the entire projection matrix at once, we apply low-rank approximation to the grouped matrix: $\mathbf{W}_{g_{j}}\approx\mathbf{L}_{g_{j}}\mathbf{R}_{g_{j}}$, where $\mathbf{L}_{g_{j}}\in\mathbb{R}^{d\times r_{g}}$ and $\mathbf{R}_{g_{j}}\in\mathbb{R}^{r_{g}\times(d_{h}\cdot s)}$. During inference, the latent representation shared across all heads in the group is computed as: $\mathbf{z}_{g_{j}}=\mathbf{x}\mathbf{L}_{g_{j}}$, and the projected outputs for individual heads are reconstructed via:
$\mathbf{y}_{j,1},\dots,\mathbf{y}_{j,s}]=\mathbf{z}_{g_{j}}\mathbf{R}_{g_{j}}.$ (4)
@@
CKA-based Head Similarity. A key question in grouped SVD is how to assign attention heads into groups so as to minimize the reconstruction error. Empirically, we observe that grouping heads with similar left singular subspaces results in lower approximation error, as these heads tend to share more common representational components and thus benefit from joint compression. To quantify head similarity, we adopt centered kernel alignment (CKA) (Kornblith et al., 2019), a robust and widely used metric for comparing representation subspaces. We compute the pairwise CKA similarity between all attention heads, yielding a symmetric similarity matrix $\mathbf{S}\in\mathbb{R}^{h\times h}$, where $h$ is the number of heads:
@@
Head Reordering. Based on the similarity matrix $\mathbf{S}$, we perform head grouping by prioritizing pairs with high mutual similarity. Specifically, we adopt a greedy strategy that iteratively selects the head pair with the highest CKA similarity and assigns them to the same group, subject to a fixed group size constraint (e.g., 4 heads per group when $h=32$). Remaining unassigned heads are then added to existing groups with available capacity, ensuring that all heads are eventually grouped. This head reordering process encourages heads with similar representational structure to share SVD decompositions, effectively reducing approximation error during grouped compression. As shown in Figure 2, we visualize the CKA similarity matrices before and after head reordering. It can be observed that, after reordering, heads assigned to adjacent positions exhibit higher mutual similarity, indicating that similar heads are more effectively grouped together. By aligning structurally coherent heads within each group, we improve both the compactness and accuracy of the resulting low-rank representation. To ensure decoding equivalence, we apply an inverse reordering to restore the original head order (see Figure 3); related theoretical analysis is given in the supplementary file.
3.3 Offline Value Calibration